OpenStudy (xmoses1):

A balloon has a volume of 10,500 liters, and the temperature is 15°C. If the temperature were -25°C, what would the volume of the balloon be? 12,194 L 9,042 L

OpenStudy (anonymous):

\[\frac{ V1 }{V 2 }=\frac{ T1 }{T2 }\] T1=15+173=288 K T2=-25+273=248 K V1=10500 L Calculate V2

OpenStudy (xmoses1):

12,194 L Is the right answer then! thanks Surjithayer :)

OpenStudy (anonymous):

i think you have calculated wrong,lower the temperature less the volume. \[\frac{ 10500 }{V2 }=\frac{ 288 }{248 },V2=\frac{ 248 }{288 }\times 10500 =?\]

OpenStudy (xmoses1):

Oh.. yeah you are right. Dang it lol oh well